how do i solve this problem using matrix?
x/6 +y/10=3
5x/16 - 3y/16=15/8
Can you also show me the cramer method. I'm lost with the fractions in the equation.

Answer by mananth(16946)

x/6 +y/10=3
LCD=30
5x +3y / 30 =3
5x+3y=30*3
5x+3y=90..............1
5x/16 - 3y/16=15/8
denominator is same
(5x-3y)/ 16 =15/8
8(5x-3y)=16*15
divide by 8
5x-3y=2*15............2
...
5 3 90
5 -3 30
D= (5*-3) - (5*3) = -30
...
Dx = (90*-3)-(30*3)=-360
..
Dy=(5*30)-(5*90)=-300
...
x= Dx/D= -360/-30= 12
y=Dy/D= -300/-30=10
...
m.ananth@hotmail.ca